OpenStudy (anonymous):

Points A and B lie on the graph of the equation y = 7x. What are the y-coordinates for each point? Point A: (−2, ) Point B: (4, ) Cora is graphing the equation . She has placed a point on the origin, and uses the slope to find 2 more points. Which points does Cora plot on her graph? Choose exactly two answers that are correct. A. (−1, −12) B. (−12, −1) C. (1, 12) D. (12, 1) What is the equation of the graphed line? Hint: Determine the slope of the line. y = x

OpenStudy (anonymous):

to find the y coordinates for the first two, plug the given x values into the given equation: y=7(-2) and y=7(4)

OpenStudy (anonymous):

so is 7,-2 and 7, 4 the answer?

OpenStudy (anonymous):

you have to multiply those numbers... I'm sorry, I should hav specified that I was just plugging the two x values in to the equation and there is an implied multiplication there... so the x value for the first one is -2, you plug that into the equation and multiply it by 7... which equals -14. so the final answer is (-2, -14)

OpenStudy (anonymous):

the answer to the second one is (4, 28)

OpenStudy (anonymous):

can you help me on any other question?

OpenStudy (anonymous):

do the multiple choice answers go with the question about Cora?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm sorry?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

does the next question have four answers to choose from? Or are those their own question?

OpenStudy (anonymous):

Ya it has multiple choice but there are 2 correct answers

OpenStudy (anonymous):

okay. The way I would do that one is by taking the answers and seeing which of them are actually on the graph

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so if you plug that in, the answer is -7. so A is not a correct answer.

OpenStudy (anonymous):

Do you understand what I did?

OpenStudy (anonymous):

sorry im back

OpenStudy (anonymous):

ok ya but there is a second answer 2

OpenStudy (anonymous):

yeah. do you think you could try it yourself?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

I just want to know if I am actually helping at all, I dont' mean to be rude....

OpenStudy (anonymous):

*don't

OpenStudy (anonymous):

and okay

OpenStudy (anonymous):

but it might be a bit... I'm also doing my own math homework, so....

OpenStudy (anonymous):

No your the only one whose been helping me on this and its a ton of help

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay! Thank you!

OpenStudy (anonymous):

Wait a second, I'm sorry, but are the first and the second question related or separate? (is the equation in the first given in the second?)

OpenStudy (anonymous):

Seperate

OpenStudy (anonymous):

Okay I'm sorry those answers I gave aren't correct...

OpenStudy (anonymous):

wait what do you mean are the seperate?

OpenStudy (anonymous):

Thery don't give you an equation in the second question.... they only give you the four answers and that the origin is on the line?

OpenStudy (anonymous):

oh sorry this got left out the equation is y=12x sorry

OpenStudy (anonymous):

Okay that makes things a lot easier

OpenStudy (anonymous):

If you dont mind I do need help with another question to thats not up there!?

OpenStudy (anonymous):

The table shows the amount of money a plumber charges for different numbers of hours that he works. The plumber graphed this information. What is the slope of the line that represents this situation? Hours worked 2 4 7 9 Amount charged ($) 96 192 336 432

OpenStudy (anonymous):

so, just like before, plug the x values into the given equation: y=12x A: (-1,-12) so y=12*(-1)=-12 B: (-12, -1) so y=12*(-12)=-144 C and D have the same answers as A and B respectively, except there aren't any negatives and Im sorry but I have to get off now so I can't help you with that last problem Sorry!